OpenStudy (anonymous):

What are the solutions? 1/2x^2 + 2x + 3 = 0

OpenStudy (tkhunny):

Have you considered the Quadratic Formula?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x ^{2} + 2x + 3 = 0\]

OpenStudy (anonymous):

@tkhunny yes. I have but it is looking for you to factor it in simplest form then set that equal to zero and find the answer that way I just don't know how with the 1/2 in front of x^2

OpenStudy (tkhunny):

Multiply the whole thing by '2' and do away with the ½, if that's a problem. Still, no matter what it says, the QuadraticFormula will guide you. It will even tell you how to factor it if that is what you seek.

OpenStudy (anonymous):

x^2+4x+6=0

OpenStudy (anonymous):

and later use quadratic formula

OpenStudy (anonymous):

THANKS @myko & @tkhunny Understand now :)